Consider the following numerical example of the simple Keynesian model with no government spending, taxes, or a foreign sector (all figures in R millions): C = 100 + 0,9Y I = 50 Answer the following questions.
(a) What is the value of the marginal propensity to consume (MPC) in this model? 2 (b) Use a graph to illustrate the equilibrium level of output.5 (c) Calculate the equilibrium level of output.4 (d) In equilibrium, what is the value of consumption spending? Use this number to verify that the sum of C and I in equilibrium equals the value for equilibrium output you obtained above.5 (e) What is the value of the multiplier in this economy?3 (f) Suppose the level of output that creates full employment in the economy is 1 800. Using the multiplier, determine the level of investment spending that would create full employment in this economy.5
1
Expert's answer
2020-12-14T05:12:56-0500

a)MPC=0.9

b) S=I





c)Y=C+I

Y=100 + 0.9Y+50

Y-0.9Y=150

0.1Y=150

Y=1500

d)C = 100 + 0.9Y

"C=100+0.9\times1500=1450"

e)"m=\frac{1}{1-MPC}=\frac{1}{1-0.9}=10"

f)"m=\frac{\Delta Y}{\Delta I}"

m=10

"10=\frac{1800-1500}{\Delta I}=\frac{300}{\Delta I}"

"\Delta I=30"

level of investment spending that would create full employment in this economy:

50+30=80